OpenStudy (anonymous):

I NEED HELP!!!! PLEASE!!! 3(x+1)=-2(x-1)-4 1 -1 -5 -25

OpenStudy (anonymous):

@kay01 -1

OpenStudy (anonymous):

Thank you .. Can you show me how you got that?

OpenStudy (anonymous):

@cutelilcigarette

OpenStudy (anonymous):

@kay01 heres your equation -> 3(x+1) = -2x (x-1) -4 step 1: simplify both sides of the equation 3x + 3 = -2x - 2 step 2: add 2x to both sides 3x + 3 + 2x = -2x - 2 + 2x 5x + 3 = -2 step 3: subtract 3 from both sides 5x + 3 - 3 = - 2 - 3 5x = - 5 step 4: divide both sides by 5 5x/5 = -5/x your answer -> x = -1